A heat engine working between two temperatures could theoretically convert one-eighth of the heat supplied to the work. If the lowered temperature is reduced by 95°C , the theoretical efficiency would be doubled. Find the source and sink temperature
1
Expert's answer
2020-08-24T13:10:40-0400

Efficiency "η = \frac{(T_{1} – T_{2})}{T_{1}} = \frac{1}{8}" (1)

T1 is the source temperature

T2 is the sink temperature

"η = \frac{T_{1} – (T_{2} - 95)}{T_{1}} = 2\times \frac{1}{8} = \frac{1}{4}"

4T1 – 4T2 + 380 = T1

3T1 – 4T2 + 380 = 0

Substituting the value of T2 from equation (1)

8T1 – 8T2 = T1

7T1 – 8T2 = 0

7T1 = 8T2

"\frac{7}{2}" T1 = 4T2


3T1"\frac{7}{2}" T1 + 380 = 0

"\frac{1}{2}" T1 = 380

T1 = 760 K or 487 °C

T2 = 665 K or 392 °C

Answer: source temperature 487 °C, sink temperature 392 °C